OpenStudy (jasmineetta):

ALGEBRA 1 Choose the equation below that represents the line passing through the point (−2, −3) with a slope of −6. A.) y + 3 = −6(x + 2) B.) y + 2 = 6(x + 3) C.) y − 3 = −6(x − 2) D.) y − 2 = 6(x − 3)

OpenStudy (bibby):

do you know what point-slope form looks like?

OpenStudy (jasmineetta):

y = mx + b... right??

OpenStudy (bibby):

\(y-y_1=m(x-x_1)\) here, \((x_1,y_1)\) represents a point the line goes through and m represents the slope

OpenStudy (jasmineetta):

Oh Okay

OpenStudy (bibby):

so just plug in the information and you should see your answer choice

OpenStudy (jasmineetta):

y − 2 = 6(x − 3) so it's D

OpenStudy (bibby):

close, but there are some adjustments to be made, I'm gonna try and color code this

OpenStudy (jasmineetta):

okay :)

OpenStudy (bibby):

\(y-\color{red}{y_1}=\color{green}{m}(x-\color{purple}{x_1})\\(\color{purple}{x_1},\color{red}{y_1})=(\color{purple}{-2},\color{red}{-3})\\\color{green}{m}=\color{green}{-6} \\ y-\color{red}({-3})=\color{green}{-6}(x-\color{purple}({-2}))\)

OpenStudy (jasmineetta):

That makes sense!! THANK YOU BIBBY!! Lol & i'ma send you a friend request on PA

OpenStudy (bibby):

slide into those dms baby girl

OpenStudy (jasmineetta):

hahaha XD